# Accelerating Multimodal Large Language Models via Dynamic Visual-Token Exit and the Empirical Findings

### v1

@@ 3.2 Optimization
The target of DyVTE is to correctly predict the exit layer of visual tokens without affecting its final outputs. In this case, its objective can be defined by
@@
where $\mathbf{\theta}_{h}$ denotes the weights of hyper-networks in DyVTE. $d(\cdot,\cdot)$ is an distance functions [31]. And $P^{\prime}_{l}$ and $P$ are the predictions of an MLLM with and without early visual token exit, respectively.
To directly optimize Eq.4, a viable way is to merge predictions of hyper-networks with visual tokens of MLLMs, e.g., attention masks based on Gumbel softamx, thus using the next token prediction loss to indirectly update DyVTE. Although the MLLM can be fixed, but the complete computation of the whole model’s gradients are still required, which is often inefficient and expensive.
Considering the target of DyVTE is to find out the optimal step for visual-token exit, we can actually train the hyper-networks independent to the gradient back-propagation of MLLMs, i.e., directly proper supervision.
Concretely, we can compare the discrete outputs of the MLLM with and without DyVTE, e.g., the answer strings $A$. If the answers are exactly the same, we can give a positive feedback to hyper-networks, and vice verse. By comparing it with the default output $A$, we can judge that whether the current input should be exited at $i$-th layer:
$\mathbf{y}=\begin{cases}1,&A^{\prime}_{l}=A,\\ 0,&A^{\prime}_{l}\neq A.\end{cases}$ (5)
Here, $A^{\prime}_{l}$ denotes the answer predicted with visual-token exit at the $i$-th layer. Besides, to make this supervision more robust, we also consider the prediction uncertainty, making the model behaviors akin more to the case without exit:
$\mathbf{y}=\begin{cases}1,&A^{\prime}_{l}=A\wedge\rho^{\prime}_{c}<\tau,\\ 0,&otherwise.\end{cases}$ (6)
Here, $\rho^{\prime}_{c}$ denotes the prediction uncertainty, represented by cross-entropy, and $\tau$ denotes the threshold, which is the default uncertainty $\rho_{c}$ multiplied by a scale value.
@@
In practice, we randomly sample exit layers at each training step, and give the weak labels according to Eq.6, based on which the hyper-networks are optimized by Eq.7.
@@ 4 Experiment
Figure 3: Distributions of averaged attention scores of image self-attention, cross-attention, and text self-attention. We visualize the mean and variance of the attention weight of each part on the LLaVA-1.5 7B, LLaVA-1.5 13B, Eagle-X5 7B and InternVL 7B. From these distributions, we can summarize three main stages of MLLMs as introduced in the main paper, which are then marked by red lines.
Figure 4: The relationship between manual visual-token exit and performance on LLaVA-1.5 7B and 13B. We show the results of removing all visual tokens at 00-th, $8$-th, $16$-th, $24$-th and $32$-th layers. “Baseline” represents the original performance of the MLLM. After a certain layer, the removal of all visual tokens barely impedes performance.
@@
InternVL 7B [11] 59.2 16.0 1525.1 15.5 64.6 16.2 86.4 15.4 31.2 15.4
Intern-DyVTE 7B 59.1 (-0.2%) 11.9 (-25.6%) 1474.1 (-3.3%) 10.9 (-29.7%) 64.4 (-0.3%) 12.0 (-25.9%) 81.3 (-5.9%) 10.9 (-29.2%) 29.5 (-5.4%) 13.0 (-15.6%)
LLaVA-1.5 7B [40] 58.6 9.2 1510.7 8.9 64.3 9.6 85.9 8.8 30.5 8.7
LLaVA-DyVTE 7B 58.6 (0.0%) 5.0 (-45.7%) 1491.4 (-1.3%) 6.3 (-29.2%) 64.7 (+0.6%) 5.4 (-43.8%) 81.6 (-5.0%) 4.1 (-53.4%) 31.9 (+4.6%) 6.3 (-27.6%)
LLaVA-1.5 13B [40] 61.6 17.6 1531.3 16.9 67.7 18.3 85.9 16.8 36.1 16.7
@@
Eagle-X5 7B [45] 64.9 27.8 83.4 27.8 71.2 29.5 69.8 29.2 72.3 28.6
Eagle-DyVTE 7B 62.4(-3.9%) 21.7 (-21.9%) 82.6 (-1.0%) 21.6 (22.3%) 70.2 (-1.4%) 24.5(-16.8%) 71.7 (+2.7%) 23.5 (-24.3%) 71.7 (-0.8%) 22.8 (-20.3%)
VILA 7B [37] 63.1 8.8 80.3 8.8 62.6 9.5 69.5 9.8 68.8 9.2
@@
InternVL 7B [11] 62.9 15.4 79.3 15.4 57.0 16.1 66.2 16.4 66.4 15.8
Intern-DyVTE 7B 61.3 (-2.5%) 11.8 (-23.4%) 77.6 (-2.1%) 11.7 (-24.0%) 55.8 (-2.1%) 13.5 (-16.1%) 66.2 (0.0%) 12.1 (-26.2%) 65.2 (-1.8%) 12.3 (-22.2%)
LLaVA-1.5 7B [40] 62.0 8.8 78.5 8.8 58.2 9.5 69.4 9.8 67.0 9.2
@@
Figure 6: Statistics of the exiting layers decided by DyVTE on LLaVA-1.5 7B, LLaVA-1.5 13B and InternVL 7B. The horizontal axis represents the exit layer, and the vertical axis represents the proportion. The exit time by is different for different tasks, but similar for the MLLMs with the same sizes.
4.1 Datasets and Metrics
The benchmarks used in this paper consist of four conventional vision-language (VL) benchmarks and five newly introduced MLLM benchmarks. The traditional VL benchmarks include VQAv2 [26], GQA [27], ScienceQA [42], and TextVQA [46]. The MLLM benchmarks comprise POPE [35], MME [24], MMB [41], SEED [32] and MM-Vet [60]. Unlike general VL evaluations, these benchmarks are designed to assess MLLMs from diverse views, such as fine-grained reasoning [60] and visual hallucination [35].
@@ 4.1 Datasets and Metrics
Mean Visual Last Visual Mean Text Last Text Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$
@@
LLaVA[40] 69.4 9.8 30.5 8.7 58.6 9.2 64.3 9.6 55.7 9.3
ToMe [4] 69.6 (+0.3%) 5.9 (-39.8%) 30.6 (+0.3%) 4.9 (-43.7%) 57.8 (-1.4%) 5.5 (-40.2%) 63.7 (-0.9%) 5.7 (-40.6%) 55.4 (-0.5%) 5.5 (-40.9%)
FastV [8] 69.0 (-0.6%) 6.2 (-36.7%) 31.3 (+2.6%) 5.2 (-35.8%) 58.2 (-0.7%) 5.8 (-37.0%) 64.4 (+0.2%) 6.0 (-37.5%) 55.7 (0.0%) 5.8 (-37.6%)
@@
4.2 Implementation Details
We apply DyVTE to three popular MLLMs, namely Eagle-X5 7B [45], VILA 7B [37], InternVL [11], LLaVA-1.5 7B [40] and LLaVA-1.5 13B [40]. Hyper-networks in the different layers do not share parameters. The scale $\alpha$ is set to $1.03$ to filter out samples with large deviations in the generated distribution. For all MLLMs, the hidden dimension of the hyper-network is set to 2,048. During training, we freeze the original MLLMs, and the training rate of hyper-networks is set to $4\times 10^{-5}$. And we randomly sample 1% from the training set of the 665k instruction data of LLaVA-1.5 [40] to train hyper-networks. The training epoch is set to 1. More details can refer to our code project.
4.3 Quantitative Analysis
4.3.1 Attention behaviors and visual redundancy
We first quantitatively investigate the attention behaviors of a set of MLLMs, and show that how the three main stages mentioned are summarized. Afterwards, we demonstrate the visual redundancy in the multimodal reasoning stage.
Attention behaviors of MLLMs. We first visualize the attention patterns of 4 MLLMs in Fig.3, including visual self-attention, visual-text cross-attention and text self-attention. Here, the values in these distributions are the averaged attention scores on the LLaVA-split [40], which can reflect the intensity of different attention modes. From these plots, we can first observe that four MLLMs share similar patterns for different types of attention. At the shallow layers, all MLLMs exhibit obvious cross-attention interactions between visual and text tokens, suggesting the intensive interactions between two modalities. This corresponds to the early fusion stage we termed above, which refers to the quick exchange of multimodal information. But this activity will quickly decline, and we can observe that the self-attention modeling within visual and text tokens becomes more active, e.g., between two red lines. In this case, we can assume that MLLM focuses starts the intra-modality modeling stage. After that, the cross-attention gradually resumes and shows a certain degree of vibration. At the same time, the other attentions present relatively irregular trends. Here, we term this phase as the multimodal reasoning stage. Notably, we can find these patterns exists in four MLLMs of different families and scales.
The impact of dropping all visual tokens. Next, we estimate the visual redundancy of MLLMs given their inference patterns discussed above. In Fig.4, we present the results of removing visual tokens on LLaVA-1.5 7B and 13B for four benchmarks with different exiting strategies. From these plots, we can first observe that visual tokens are not always required throughout the entire process, especially the last stage. For instance, on GQA, removing all visual tokens at $24$-th layer only has $0.5\%$ and $0.6\%$ performance drops for LLaVA-1.5 7B and 13B, receptively. Another finding is that the exit time of visual tokens is different for different tasks or examples. For instance, the optimal exit layer of SQA is earlier than that of TextVQA, i.e., 16 v.s. 25, suggesting the need of DyVTER. More importantly, we can find that the correct removal of visual tokens has little impact on the self-attention patterns of text tokens. As shown in Fig.5, the text self-attention still presents similar distributions after using DyVTE, regardless of the self-modeling of visual tokens. The only difference lies in the slight change of attention intensity due to the shorter token sequence. Overall, these results well validate that visual tokens barely make contributions after a certain phase, and dynamically removing all visual tokens has little impact on the reasoning.
4.3.2 Results of DyVTE
@@
The attention entropy with visual tokens exit. In Fig.5, we visualize the entropy distributions of cross-attention matrices and text self-attention matrices of LLaVA-1.5 7B, and compare the results with and without DyVTE across four benchmarks. We first observe that the processes of multimodal reasoning are generally similar across these benchmarks. In the text self-attention matrix, which is involved in answer generation, the entropy distributions across these datasets are quite similar. Another key observation is that visual tokens are often redundant in multimodal reasoning. Specifically, after removing all visual tokens, cross-modality information transfer is no longer performed, yet the text self-attention remains largely unaffected, as discussed above. Additionally, our DyVTE method can effectively remove all visual tokens at a specific layer according to the status of multimodal reasoning. In particular, the exit determined by DyVTE only blocks cross-modality information diffusion while preserving the modeling of text tokens. Overall, these findings support our motivation that visual tokens becomes redundant before the end of inference.
Results of DyVTE on different MLLMs. In Tab.2 and Tab.2, we present the results of applying DyVTE to a set of MLLMs of different families and sizes, including Eagle [45], VILA [37], InternVL [11], and LLaVA-1.5 [40]. From the tables, we can first observe that DyVTE significantly reduces computational overhead in existing MLLMs. For example, LLaVA-DyVTE 7B achieves a 45.7% reduction in FLOPs on SEED without sacrificing performance. Additionally, the extent of FLOPs reduction varies depending on the task. On SQA with multiple-choice questions, DyVTE reduces computational overhead by up to 43.9%, while on the more detailed MM-Vet dataset, the reduction is 27.6%. Another notable observation is that DyVTE’s impact varies across different MLLMs. Specifically, VILA-DyVTE 7B reduces computational overhead by 48.3% on the MME benchmark with no performance drop, whereas LLaVA-DyVTE 7B only reduces overhead by 29.2%. When applying DyVTE to larger MLLMs, such as LLaVA-1.5 13B, we can observe a more significant efficiency improvement, i.e., LLaVA-DyVTE 13B reduces computational overhead by 55.9% on SQA, with only a 0.8% performance drop. Furthermore, DyVTE proves to be particularly effective for MLLMs with strong visual representations. For instance, when DyVTE is applied to Eagle-X5 7B, which has a visual encoder containing 13.3 TFLOPs, the performance drop is minimal, i.e. only 0.8% on traditional benchmarks, while Eagle-DyVTE 7B achieves a reduction of up to 6.2 TFLOPs in computational overhead.
The distribution of exit layers. In Fig. 6, we present the distributions of the visual token exit layers for LLaVA-DyVTE 7B and 13B. From the figure, we observe that DyVTE dynamically selects the exit layer based on the specific task requirements. For example, on the TextVQA dataset, which demands fine-grained information, the exit occurs later compared to other datasets. In contrast, on the GQA dataset, which involves open-ended word questions and simple true/false judgments, the exit layers are more evenly distributed across early and late layers. Another notable observation is that the timing of the exit layer shows similar distributions across MLLMs of different scales and architectures. For instance, both LLaVA-DyVTE 7B and 13B remove all visual tokens at the same layer early in the process on the SQA dataset. Similarly, for InternVL and LLaVA, peaks in the exit layer distributions occur at comparable layers. Overall, these results support our motivation and the proposed DyVTE.
@@
Ablation study. In Tab.3, we perform a set of experiments to analyze the impact of different representation selections on DyVTE. In this table, various tokens are used as the token status in Eq.1. As shown in Tab.3, when relying on image-only tokens, the model tends to remove all visual tokens at a later stage, specifically after the $21$-st layer. Using the “mean text” token status allows for an earlier exit from the layers. When only the “last text” token is used, DyVTE removes all visual tokens prematurely, e.g., the $17$-th layer, resulting in a more obvious performance drop of 1.0% on average. For combinations of two representations, we can observe that the “mean text + last text” is more effective. Specifically, it removes visual tokens at the about $17$-th layer, while causing only a 0.5% performance drop. On the other hand, using a combination of visual token statuses, i.e., “mean Visual + last visual”, yields results similar to using a single visual representation, with visual tokens being retained until the $21$-st layer. Overall, these results well confirm the use of text token status for dynamic visual-token exiting in MLLMs.
Figure 7: Examples on LLaVA-1.5 with DyVTE. Our DvVTE can help the MLLM answer the questions as accurately as default, while being faster, i.e., more tokens per second.
@@
Comparison with token pruning methods. In Tab.4, we compare the performance and efficiency of DyVTE with the visual token pruning methods of different principles. From the table, we can observe that DyVTE outperforms token pruning methods in both performance and efficiency. For example, when compared with ToMe on SQA, DyVTE not only maintains better performance but also reduces FLOPs by 4.1%. In comparison to FastV, the advantage of DyVTE can be more pronounced. Specifically, DyVTE improves performance while reducing computation overhead by 8.7% on the SEED benchmark. Another key observation is that DyVTE exhibits significant performance gains on more complex tasks. For instance, on the MM-Vet benchmark, which evaluates multiple capabilities including math and OCR, DyVTE improves performance by 4.3% compared to ToMe. Additionally, DyVTE demonstrates excellent compatibility with existing methods. When combined with FastV, DyVTE reduces computation overhead by 51.5%, with only a 0.7% performance drop on average. Overall, these results validate the effectiveness and efficiency of DyVTE, establishing it as a powerful tool in efficient MLLM development.

### v2

@@ 4.2 Optimization
In particular, given sufficient examples of visual-token exit, DyVTE can learn to judge the text token status at each layer of the MLLM. Thus, its accurate removals of visual tokens will greatly speed up the inference in the rest layers of the MLLM, while retaining similar predictions.
Thus the objective of DyVTE can be defined by
@@
where $\mathbf{\theta}_{h}$ denotes the weights of hyper-networks in DyVTE. $d(\cdot,\cdot)$ is KL-Divergence kullback1951information . And $P^{\prime}_{l}$ and $P$ are the predictions of an MLLM with and without early visual token exit, respectively.
To optimize Eq.3, a direct approach is to merge the predictions of hyper-networks with visual tokens of MLLMs, e.g., implementing attention masks based on Gumbel softmax jang2016categorical , thus using the next token prediction loss to indirectly update DyVTE. Although feasible, this approach still requires the computation of the model’s all gradients, which is still inefficient and expensive.
In this case, we propose an efficient training regime independent to the gradient back-propagation of MLLMs. Concretely, we can compare the discrete outputs of the MLLM with and without DyVTE, e.g., the answer strings $A$. If the answers are exactly the same, we can give a positive feedback to hyper-networks, and vice verse. By comparing it with the default output $A$, we can judge that whether the visual tokens should be exited at $l$-th layer:
[…]
Here, $A^{\prime}_{l}$ denotes the answer predicted with visual-token exit at the $l$-th layer. Besides, to make this supervision more robust, we also consider the prediction uncertainty as a regularization term, thereby , making the MLLM behaviors closer to the default inference:
[…]
Here, $\rho^{\prime}_{c}$ denotes the prediction uncertainty valued by cross-entropy and multiplied by a scaling factor, and $\tau$ denotes the threshold.
@@
Compared with the indirect optimization using next token prediction PyramidDrop , Eq.5 can provide more effective supervisions to DyVTE. We randomly sample exit layers at each training step, and compute the labels $\mathbf{y}$ according to Eq.5. Finally, the hyper-networks are optimized based loss defined in Eq.6. The overall expenditure of DyVTE training will be much cheaper than tuning MLLMs or learning new prediction layers in previous dynamic exit approaches ee-llm ; PyramidDrop .
@@ 5 Experiment
5.1 Datasets and Metrics
The benchmarks used in this paper consist of four conventional vision-language (VL) benchmarks and five newly introduced MLLM benchmarks. The traditional VL benchmarks include VQAv2 vqav2 , GQA gqa , ScienceQA sqa-i , and TextVQA textvqa . The MLLM benchmarks comprise POPE pope , MME mme , MMB mmbench , SEED seed and MM-Vet mm-vet . Different from general VL evaluation, the MLLM benchmarks focus more on the evaluations of MLLMs, such as fine-grained reasoning mm-vet and visual hallucination pope .
@@ 5.1 Datasets and Metrics
InternVL 7B 59.2 16.0 1525.1 15.5 64.6 16.2 86.4 15.4 31.2 15.4
InternVL-DyVTE 7B 59.1 (-0.2%) 11.9 (-25.6%) 1474.1 (-3.3%) 10.9 (-29.7%) 64.4 (-0.3%) 12.0 (-25.9%) 81.3 (-5.9%) 10.9 (-29.2%) 29.5 (-5.4%) 13.0 (-15.6%)
LLaVA-1.5 7B 58.6 9.2 1510.7 8.9 64.3 9.6 85.9 8.8 30.5 8.7
LLaVA-DyVTE 7B 58.6 (0.0%) 5.0 (-45.7%) 1491.4 (-1.3%) 4.3 (-51.7%) 64.7 (+0.6%) 5.4 (-43.8%) 81.6 (-5.0%) 4.1 (-53.4%) 31.9 (+4.6%) 6.3 (-27.6%)
LLaVA-1.5 13B 61.6 17.6 1531.3 16.9 67.7 18.3 85.9 16.8 36.1 16.7
@@
EAGLE-X5 7B 64.9 27.8 83.4 27.8 71.2 29.5 69.8 29.2 72.3 28.6
EAGLE-DyVTE 7B 62.4 (-3.9%) 21.7 (-21.9%) 82.6 (-1.0%) 21.6 (-22.3%) 70.2 (-1.4%) 24.5 (-16.8%) 71.7 (+2.7%) 23.5 (-24.3%) 71.7 (-0.8%) 22.8 (-20.3%)
VILA 7B 63.1 8.8 80.3 8.8 62.6 9.5 69.5 9.8 68.8 9.2
@@
InternVL 7B 62.9 15.4 79.3 15.4 57.0 16.1 66.2 16.4 66.4 15.8
InternVL-DyVTE 7B 61.3 (-2.5%) 11.8 (-23.4%) 77.6 (-2.1%) 11.7 (-24.0%) 55.8 (-2.1%) 13.5 (-16.1%) 66.2 (0.0%) 12.1 (-26.2%) 65.2 (-1.8%) 12.3 (-22.2%)
LLaVA-1.5 7B 62.0 8.8 78.5 8.8 58.2 9.5 69.4 9.8 67.0 9.2
@@
5.2 Implementation Details
We apply DyVTE to five popular MLLMs, namely EAGLE-X5 7B EAGLE , VILA 7B VILA , InternVL 7B InternVL1 , LLaVA-1.5 7B llava-1.5 and LLaVA-1.5 13B llava-1.5 . The scale for $\rho^{\prime}$ in Eq.5 is set to 1.031.031.03. For all MLLMs, the hidden dimension of the hyper-network is set to 2,048. During training, all MLLMs are frozen, and the training rate of hyper-networks is set to $4\times 10^{-5}$. And we randomly sample 1% examples of the LLaVA-665k instruction set llava-1.5 to train DyVTE for 1 epoch. More details can refer to our code project.
5.3 Quantitative Analysis
Results of DyVTE on different MLLMs. In Tab.2 and Tab.2, we report the results of applying DyVTE to a set of MLLMs of different families and sizes, including EAGLE EAGLE , VILA VILA , InternVL InternVL1 , and LLaVA-1.5 llava-1.5 . From these tables, we can first observe that DyVTE significantly reduces the computational overhead of existing MLLMs. For example, DyVTE reduces FLOPs of LLaVA 7B by 45.7% on SEED without dropping performance. Additionally, the actual reduction of FLOPs are distinct for different tasks. For instance, on SQA with multiple-choice questions, DyVTE reduces computational overhead by up to 43.9%, and on MM-Vet about granular answers, the reduction achieved is 27.6%. These results confirm our intuition that the optimal removal of visual tokens are different for different tasks. Another observation is that DyVTE’s effects vary across these MLLMs.
Table 3: Comparison of real inference budget between token pruning methods and vanilla decoding. “Acc.” denotes the accuracy.
Method SQA-I MMB
Latency $\downarrow$ Acc. $\uparrow$ Latency $\downarrow$ Acc. $\uparrow$
LLaVA 13B 0.237s 72.9 0.236s 67.7
FastV 0.171s (-27.7%) 73.1 (+0.3%) 0.174s (-26.2%) 68.6 (+1.3%)
ToMe 0.175s (-26.2%) 73.2 (+0.4%) 0.179s (-24.2%) 67.4 (-0.4%)
DyVTE 0.161s (-32.1%) 72.3 (-0.8%) 0.163s (-30.9%) 66.0 (-2.5%)
DyVTE+FastV 0.154s (-34.9%) 73.0 (+0.2%) 0.155s (-34.4%) 68.5 (+1.2%)
Specifically, DyVTE can reduce the computational overhead of VILA 7B by 48.3% on the MME benchmark with no performance drop, whereas InternVL-DyVTE 7B only achieve a reduction of 29.7%. When applying DyVTE to larger MLLMs, such as LLaVA-1.5 13B, we can observe a more significant efficiency improvement. For instance, DyVTE reduces the computational overhead of LLaVA-1.5 13B by 55.9% on SQA with only -0.8% on performance. Furthermore, DyVTE is more effective for MLLMs that uses stronger visual representations. For instance, EAGLE-X5 is a SOTA MLLM with strong and hybrid visual backbones, and it works very well with our DyVTE. DyVTE can reduce its computation by up to 6.2 TFLOPs with only 0.8% performance drop on the VL benchmarks. Overall, these results not only validate the effectiveness of our DyVTE but also confirm our motivations and findings about MLLMs.
Table 4: Ablation study of different token statuses for DyVTE. “Mean Visual” denotes the average of all visual tokens except the last one, similar with “Mean Text”. “Last Visual” and “Last Text” denotes the last visual or text token. “Exit Layer” denotes the averaged layer numbers selected to remove by DyVTE. “Acc.” denotes the accuracy. The default setting of DyVTE is the last row. The best and second best results are marked in bold and underline, receptively.
@@ 5.3 Quantitative Analysis
Mean
Visual
Last
Visual
Mean
Text
Last
Text
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
@@
LLaVA 7B 69.4 9.8 30.5 8.7 58.6 9.2 64.3 9.6 55.7 9.3
ToMe ToMe 69.6 (+0.3%) 5.9 (-39.8%) 30.6 (+0.3%) 4.9 (-43.7%) 57.8 (-1.4%) 5.5 (-40.2%) 63.7 (-0.9%) 5.7 (-40.6%) 55.4 (-0.5%) 5.5 (-40.9%)
FastV FastV 69.0 (-0.6%) 6.2 (-36.7%) 31.3 (+2.6%) 5.2 (-35.8%) 58.2 (-0.7%) 5.8 (-37.0%) 64.4 (+0.2%) 6.0 (-37.5%) 55.7 (0.0%) 5.8 (-37.6%)
@@
Inference Latency. In Tab. 3, we compare the actual inference latency of DyVTE with the token prune and the vanilla decoding methods. From the table, we can first observe that token prune methods can significantly reduce the inference latency. For instance, FastV can reduce the inference time by -27.7% and -26.2% on SQA and MMB benchmarks, receptively. Compared with these token prune methods, our DyVTE further enhances inference efficiency. Specifically, DyVTE improve the inference efficiency by 32.1% and 30.9% on SQA and MMB benchmarks with competitive performance. Moreover, the proposed DyVTE can work together with existing token prune methods and achieve better efficiency. For example, when combined with FastV, DyVTE achieves even greater reductions in inference latency, i.e., 34.9% and 34.4% on SQA and MMB benchmarks. Overall, these results validate our DyVTE can accelerate the inference.
@@
The attention entropy with visual tokens exit. In Fig.5, we visualize the entropy distributions of cross-attention matrices and text self-attention matrices of LLaVA-1.5 7B. And we compare the results with and without DyVTE across four benchmarks. We first observe that there are differences between the multimodal reasoning procedures across different benchmarks. Specifically, in SQA and TextVQA, the changes in entropy during text modeling are more obvious than those in GQA and MM-Vet. This phenomenon fully illustrates the necessity of adopting a dynamic approach to exit visual tokens. Another observation is that DyVTE can effectively remove all visual tokens at a specific layer according to the status of multimodal reasoning. Specifically, after removing all visual tokens by DyVTE, the entropy of text self-attention remains unaffected. In this way, the exiting determined by DyVTE only blocks visual information’s cross-modality propagation while preserving text tokens’ modeling. Overall, these findings validate that visual tokens are not always necessary in the reasoning process, and also confirm the effectiveness of the proposed DyVTE.
@@
Ablation study. In Tab.4, we perform a set of experiments to analyze the effectiveness of different token status for DyVTE. In this table, various tokens are used as the token statuses in Eq.1. As shown in Tab.4, when using only image-related tokens, the model tends to remove all visual tokens at a later stage. Specifically, the average exit layer are 21.421.421.4 and 21.221.221.2. Using the “Mean Text” token leads to an earlier exit from the layers, but still retains the computation redundancy. When only the “Last Text” token is used, DyVTE removes all visual tokens too early, i.e., at the 171717-th layer, resulting in an obvious performance drop, i.e., performance drops by 1.0% on average. For combinations of two representations, we can observe that the “Mean Text + Last Text” is the most effective way. Specifically, it removes visual tokens at about 171717-th layer, while causing only a 0.5% performance drop. On the other hand, using a combination of visual token statuses, i.e., “Mean Visual + Last Visual”, yields results similar to using a single visual representation. In this way, the visual tokens are retained until the 212121-st layer. Overall, these results confirm that the key to judging whether visual tokens have a necessary impact on the prediction lies in the status of text tokens.
@@
Comparison with token pruning methods. In Tab.5, we compare the performance and efficiency of DyVTE with the visual token pruning methods on LLaVA-1.5 7B. From the table, we can observe that DyVTE outperforms token pruning methods in both performance and efficiency. For example, when compared with ToMe on SQA, DyVTE not only maintains the performance but also further reduces FLOPs by 4.1%. When compared with FastV, the advantage of DyVTE is more significant. Specifically, DyVTE improves performance by 0.4% while reducing computation overhead by 8.7% on the SEED. Another observation is that DyVTE can significantly improve performance in complex tasks. For example, in the MM-Vet benchmark, which evaluates multiple functions including mathematics and OCR, DyVTE outperforms ToMe by 4.3Additionally, DyVTE shows excellent compatibility with existing methods. When combined with FastV, DyVTE reduces the computational overhead by 51.5% while only decreasing the performance by 0.7% on average. Overall, these results validate the effectiveness and efficiency of DyVTE.